Displacement vectors A, B, and C add up to a total of zero. Vector A has a magnitude of 1550 m and a direction of 23.4° north of east. Vector B has a direction of 41.0° east of south, and vector C has a direction of 35.9° north of west. Find the magnitudes of vector B and vector C.

.
Answer #1

Lets assume East direction as x-axis and North direction as y-axis. Then we can write given vectors in vector notation as

\vec{A}= Acos 23.4^\circ\hat{i}+ Asin 23.4^\circ\hat{j}(North of east angle will be angle of vector w.r.t. x-axis)

where A= magnitude of vector A,

  \vec{A}= 1550cos 23.4^\circ\hat{i}+ 1550sin 23.4^\circ\hat{j}

  \vec{A}= 1422.52\hat{i}+ 615.58\hat{j}

Let magnitude of vectors B and C be "B" and "C" only, then writing these vectors also in vector notation

\vec{B}= Bcos( 270^\circ+41^\circ)\hat{i}+ Bsin( 270^\circ+41^\circ)\hat{j}  (East of south angle will be angle of vector w.r.t. -y axis, then writing angle counterclockwise w.r.t. +x-axis in vector notation)

  \vec{B}= 0.656B\hat{i}-0.754B\hat{j}

Similarly writing vector C,

\vec{C}= Ccos( 180^\circ-35.9^\circ)\hat{i}+ Bsin( 180^\circ-35.9^\circ)\hat{j}  (North of west angle will be angle of vector w.r.t. -x axis, then writing angle counterclockwise w.r.t. +x-axis in vector notation)

  \vec{C}= -0.81C\hat{i}+0.586C\hat{j}

Then as given in question ,

\vec{A}+\vec{B}+\vec{C}=0

1422.52\hat{i}+615.58\hat{j}+0.656B\hat{i}-0.754B\hat{j} -0.81C\hat{i}+0.586C\hat{j}=0

equating values of same unit vector to 0,

  1422.52+0.656B-0.81C=0

C=1756.19 +0.809B...................(1)

Also,

615.58-0.754B+0.586C=0

using equation 1 in above,

  615.58-0.754B+0.586(1756.19+0.809B)=0

615.58-0.754B+1029.13+0.474B=0

  0.28B=1644.71

  B= 5873.96 m(ANS)

using above value in equation 1,

  C=1756.19 +0.809(5872.96)= 6508.223m(ANS)